HUBBLE SPACE TELESCOPE DAILY REPORT #879 PERIOD COVERED: 0800 EST Daily Status Report as of 124/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4105 (HST Medium-Deep Survey: Cycle 1 High Latitude) The Wide field Camera was used to make one high latitude medium-deep sky survey observation taken in parallel with the prime science observation. The observation was received and displayed in the OSS. The observation was reported to look okay. 1.2 Completed FGS 2861 (Linking HIPPARCOS to the Extragalactic Reference Frame) FGS #3 was used to make astrometric observations of extragalactic object NEWEGOB-35 and a group of nearby reference stars. The astrometric observations occurred mostly during LOS. There was a few minutes of telemetry near the middle of the observations. No problems were seen. 1.3 Completed PC 4104 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) The Planetary Camera was used to make two internal flat observations to check for new forms of contamination, and the possible return of quantum efficiency hysteresis. Both observations were received and displayed. They were reported to look okay. 1.4 Completed FOS/RD 4791 (Binaries in Globular Clusters (FOC 37) -- Spectroscopy) The Faint Object Spectrograph was used to make observations of stars in globular cluster NGC-6712. The binary search and the peakup were successful. The next two observations were received and displayed for the principal investigator. Hard copies were also made. No anomalies were noted. 1.5 Completed FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) The Faint Object Spectrograph was used to make spectral observations of the planetary nebulae LMC-SMP-8 and LMC-SMP-35 which are both located in the Large Magellanic Cloud. The first set of five spectra was reported to look okay, and the second set of three spectral observations were reported to have low signal and to be noisy. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed FOS/RD 4062 (Stellar and Gas dynamics in Normal Galaxies Augmentation: Cycle 2) The Faint object Spectrograph ware used to make spectral observations of nearby galaxy NGC-334 (aka, M-31) and also a set of wavelength calibration images. The binary target acquisition and peakup were successful. The spectral observation executed with no problems. The wave length calibration observation also executed and remains on the science tape recorder. 1.8 Completed FOC/96 3801 (Search for QSOs Suitable for Subsequent Observations of He II) The Faint Object Camera (f/96) was used to make observations of quasar 2054-355 to measure its ultraviolet flux levels. This is to find out if it is suitable for future observations using the FOS and GHRS. The observations were executed as scheduled, and no problems were reported. 1.9 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies With FOC/96) The Faint Object Camera was used to make snapshot images of nearby galaxy ESO-1153-1. These were ten minute exposures using filter F220W. The OSS monitored the observations and assessed the data quality. No problems were reported. 1.10 Completed FOS/BL 4054 (Probing the Nuclear Regions of Galaxy NGC-5548) The Faint Object Spectrograph (Blue side) was used to make observations of the galaxy NGC-5548 to study the time variability of its continuum, and emission lines in Helium II and Carbon IV. The observations were executed as scheduled, and no problems were reported. 1.11 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Both the Wide Field camera and the Planetary Camera were used to make internal flat observations. This was to check for new forms of contamination (i.e., measles and daisies), and the return fo quantum efficiency hysteresis. The OSS assessed the data quality, and no problems were reported. 1.12 Completed WFC 3917 (HST Medium-Deep Survey: Cycle 2 Part 1) The Wide Field Camera was used to make high latitude mediun-deep sky survey observations (Cycle 2 Part 1) taken in parallel with the prime science observations. The OSS monitored the observations, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 9 Successful : 9 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 4833 (Crane, Augmentation Cycle 3 High Spectrum of of the Jet in 3C264) 3.2 Complete WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) 3.3 Complete FGS 4775 (Exposure time Calibration in High Background Fields) 3.4 Complete FOC/96 4846 (High Resolution Observations of Cataclysmic Variables: Cycle 3) 3.5 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.6 Complete FOC/96 3801 (Search for QSOs Suitable for Subsequent Observations of He II) 3.7 Complete FOS/BL 4054 (Probing Nuclear Regions of the Seyfert Galaxy NGC-5548) 3.8 Complete WFC 3917 (HST Medium-Deep Survey: Cycle 2 Part 1) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-